# Toward Global Convergence of Gradient EM for Over-Parameterized Gaussian Mixture Models

### v1

@@ 4.3 Proof ideas for Theorem 7
For any $\bm{\mu}$ satisfying $\|\mu_{1}\|\leq\sqrt{d},\|\mu_{2}\|,\|\mu_{3}\|,\ldots,\|\mu_{n}\|\geq 10\sqrt{d}$, the gradient of $\mathcal{L}$ at $\bm{\mu}$ can be upper bounded as
$\|\nabla_{\mu_{i}}\mathcal{L}(\bm{\mu})\|\leq 2\|\mu_{1}\|+2\exp(-d)\sum_{i\neq 1}\|\mu_{i}\|,\forall i\in[n].$
@@ 5 Experiments
In this section we include a few simulation results of gradient EM verifying our theoretical arguments. We choose the experimental setting of $d=5,\eta=0.7$. We use $n=2,5,10$ Gaussian mixtures to learn data generated from one single ground truth Gaussian distribution $\mathcal{N}(\mu^{*},I_{d})$, respectively. Since a closed form expression of the population gradient is intractable, we approximate the gradient step via Monte Carlo method, with sample size $3.5\times 10^{5}$. The mixing weights of student GMM are randomly sampled from a standard Dirichlet distribution and set as fixed during gradient EM update. The covariances of all component Gaussians are set as the identity matrix. We recorded the convergence of likelihood function $\mathcal{L}$ (estimated also by Monte Carlo method on fresh samples each iteration) and parametric distance $\sum_{i\in[n]}\pi_{i}\|\mu_{i}-\mu^{*}\|^{2}$ along gradient EM trajectory. The results are reported in Figure 1 (left and middle panel). Simulation outcome indicates that both the likelihood $\mathcal{L}$ and the parametric distance converges sublinearly, verifying our theoretical arguments.
To verify our negative result Theorem 7, we consider the bad initialization point $\bm{\mu}(0)$ described in Theorem 7  [^2] To prevent numerical underflow issues, we change the constant $12$ in $\bm{\mu}(0)$ to $2$. and plot the gradient norm at $\bm{\mu}(0)$ w.r.t. different dimension $d$ in Figure 1 (right panel). Simulation result shows that the gradient norm $\|\nabla\mathcal{L}(\bm{\mu}(0))\|$ at $\bm{\mu}(0)$ decreases exponentially in dimension $d$, verifying the existence of bad initialization regions.
Figure 1: Left: Sublinear convergence of the likelihood loss $\mathcal{L}$. Middle: Sublinear convergence of the parametric distance $\sum_{i\in[n]}\pi_{i}\|\mu_{i}-\mu^{*}\|^{2}$ between student GMM and the ground truth. Right: Gradient norm $\|\nabla\mathcal{L}(\bm{\mu}(0))\|$ in the counter-example in Theorem 7 decreases exponentially fast w.r.t. dimension $d$.
@@ Appendix A Missing Proofs and Auxiliary lemmas
For any fixed $x\in\mathbf{R}^{d},x\neq 0$ and any $\bm{\mu}$ we have
$\int_{t=-1}^{1}\psi_{i}(tx|\bm{\mu})\psi_{j}(tx|\bm{\mu})\mathrm{d}t\geq\frac{1}{2\mu_{\max}\|x\|}\pi_{i}\pi_{j}\exp\left(-4\mu_{\max}^{2}\right)\left(1-\exp\left(-4\mu_{\max}\|x\|\right)\right).$
@@
$\begin{split}\int_{t=-1}^{1}\psi_{i}(tx)\psi_{j}(tx)\mathrm{d}t&\geq\int_{t=-1}^{1}\pi_{i}\pi_{j}\exp\left(-4\mu_{\max}(\|tx\|+\mu_{\max})\right)\mathrm{d}t\\ &=\pi_{i}\pi_{j}\exp\left(-4\mu_{\max}^{2}\right)\cdot 2\int_{t=0}^{1}\exp\left(-4\mu_{\max}\|x\|t\right)\mathrm{d}t\\ &=\frac{1}{2\mu_{\max}\|x\|}\pi_{i}\pi_{j}\exp\left(-4\mu_{\max}^{2}\right)\left(1-\exp\left(-4\mu_{\max}\|x\|\right)\right).\end{split}$ (13)
@@ B.1 Proofs for global convergence analysis
$\left\langle\nabla_{\bm{\mu}}Q(\bm{\mu}|\bm{\mu}),\bm{\mu}\right\rangle=\mathbf{E}_{x}[\|\tilde{\bm{\psi}}_{\bm{\mu}}(x)\|^{2}]\geq\Omega\left(\frac{\exp\left(-8\mu_{\max}^{2}\right)\pi_{\min}^{2}}{d(1+\mu_{\max}{\sqrt{d}})^{2}}\mu_{\max}^{4}\right).$
@@
$\begin{split}\mathbf{E}_{x}\left[\|\tilde{\bm{\psi}}_{\bm{\mu}}(x)\|^{2}\right]&\geq\frac{\exp\left(-8\mu_{\max}^{2}\right)}{40000d(1+2\mu_{\max}{\sqrt{d}})^{2}}\left(\sum_{i,j\in[n]}\pi_{i}\pi_{j}\|\mu_{i}-\mu_{j}\|^{2}\right)^{2}\\ &\geq\frac{\exp\left(-8\mu_{\max}^{2}\right)}{40000d(1+2\mu_{\max}{\sqrt{d}})^{2}}\left(\frac{\pi_{\min}}{8}\mu_{\max}^{2}\right)^{2}\\ &=\frac{\exp\left(-8\mu_{\max}^{2}\right)\pi_{\min}^{2}}{2560000d(1+2\mu_{\max}{\sqrt{d}})^{2}}\mu_{\max}^{4}.\end{split}$
Case2. For $\forall k\in[n]$, $\|\mu_{{i_{\max}}}-\mu_{k}\|<\frac{\mu_{\max}}{2}$. Then by Lemma 21 we have $\mathbf{E}_{x}\left[\|\tilde{\bm{\psi}}_{\bm{\mu}}(x)\|^{2}\right]\geq\frac{1}{4}\mu_{\max}^{2}\geq\Omega(\exp(-8\mu_{\max}^{2})\mu_{\max}^{4})\geq\Omega\left(\frac{\exp\left(-8\mu_{\max}^{2}\right)\pi_{\min}^{2}}{d(1+\mu_{\max}{\sqrt{d}})^{2}}\mu_{\max}^{4}\right),$ (since $e^{-x}x\leq 1,\forall x$). ∎
@@
$\mathbf{E}_{x}\left[\|\tilde{\bm{\psi}}_{\bm{\mu}}(x)\|^{2}\right]\geq\frac{\exp\left(-8\mu_{\max}^{2}\right)}{40000d(1+2\mu_{\max}{\sqrt{d}})^{2}}\left(\sum_{i,j\in[n]}\pi_{i}\pi_{j}\|\mu_{i}-\mu_{j}\|^{2}\right)^{2}.$
@@
$\begin{split}&\quad\mathbf{E}_{x}\left[\|\tilde{\bm{\psi}}_{\bm{\mu}}(x)\|^{2}\right]\\ &\geq\frac{1}{16}\mathbf{E}_{x}\left[\left(\int_{t=-1}^{1}\|x\|\sum_{i,j\in[n]}\psi_{i}(tx)\psi_{j}(tx)\langle\mu_{i}-\mu_{j},\overline{x}\rangle^{2}\mathrm{d}t\right)^{2}\right]\\ &=\frac{1}{16}\mathbf{E}_{x}\left[\left(\|x\|\sum_{i,j\in[n]}\langle\mu_{i}-\mu_{j},\overline{x}\rangle^{2}\int_{t=-1}^{1}\psi_{i}(tx)\psi_{j}(tx)\mathrm{d}t\right)^{2}\right]\\ &\geq\frac{1}{16}\mathbf{E}_{x}\left[\left(\|x\|\sum_{i,j\in[n]}\langle\mu_{i}-\mu_{j},\overline{x}\rangle^{2}\frac{1}{2\mu_{\max}\|x\|}\pi_{i}\pi_{j}\exp\left(-4\mu_{\max}^{2}\right)\left(1-\exp\left(-4\mu_{\max}\|x\|\right)\right)\right)^{2}\right]\\ &=\frac{\exp\left(-8\mu_{\max}^{2}\right)}{64}\mathbf{E}_{x}\left[\left(\sum_{i,j\in[n]}\pi_{i}\pi_{j}\langle\mu_{i}-\mu_{j},\overline{x}\rangle^{2}\frac{1-\exp\left(-4\mu_{\max}\|x\|\right)}{\mu_{\max}}\right)^{2}\right]\\ &\geq\frac{\exp\left(-8\mu_{\max}^{2}\right)}{64}\left(\sum_{i,j\in[n]}\pi_{i}\pi_{j}\mathbf{E}_{x}\left[\langle\mu_{i}-\mu_{j},\overline{x}\rangle^{2}\frac{1-\exp\left(-4\mu_{\max}\|x\|\right)}{\mu_{\max}}\right]\right)^{2}\end{split}$ (24)
@@
$\begin{split}\mathbf{E}_{x}\left[\|\tilde{\bm{\psi}}_{\bm{\mu}}(x)\|^{2}\right]&\geq\frac{\exp\left(-8\mu_{\max}^{2}\right)}{64}\left(\sum_{i,j\in[n]}\pi_{i}\pi_{j}\mathbf{E}_{x}\left[\langle\mu_{i}-\mu_{j},\overline{x}\rangle^{2}\frac{1-\exp\left(-4\mu_{\max}\|x\|\right)}{\mu_{\max}}\right]\right)^{2}\\ &\geq\frac{\exp\left(-8\mu_{\max}^{2}\right)}{64}\left(\sum_{i,j\in[n]}\pi_{i}\pi_{j}\frac{1-\exp\left(-2\mu_{\max}{\sqrt{d}}\right)}{50d\mu_{\max}}\|\mu_{i}-\mu_{j}\|^{2}\right)^{2}\\ &\geq\frac{\exp\left(-8\mu_{\max}^{2}\right)}{64}\left(\sum_{i,j\in[n]}\pi_{i}\pi_{j}\frac{1-\frac{1}{1+2\mu_{\max}{\sqrt{d}}}}{50d\mu_{\max}}\|\mu_{i}-\mu_{j}\|^{2}\right)^{2}\\ &=\frac{\exp\left(-8\mu_{\max}^{2}\right)}{40000d(1+2\mu_{\max}{\sqrt{d}})^{2}}\left(\sum_{i,j\in[n]}\pi_{i}\pi_{j}\|\mu_{i}-\mu_{j}\|^{2}\right)^{2}\end{split}$ (28)
@@
Consider training a student $n$-component GMM initialized from $\bm{\mu}(0)=(\mu_{1}(0)^{\top},\ldots,\mu_{n}(0)^{\top})^{\top}$ to learn a single-component ground truth GMM $\mathcal{N}(0,I_{d})$ with population gradient EM algorithm. If the step size satisfies $\eta\leq O\left(\frac{\exp\left(-8n\mu_{\max}^{2}(0)\right)\pi_{\min}^{2}}{n^{2}d^{2}(\frac{1}{\mu_{\max}(0)}+\mu_{\max}(0))^{2}}\right)$, then gradient EM converges globally with rate
@@
where constant $\gamma=\Omega\left(\frac{\eta\exp\left(-16n\mu_{\max}^{2}(0)\right)\pi_{\min}^{4}}{n^{2}d^{2}(1+\mu_{\max}(0){\sqrt{dn}})^{4}}\right)\in\mathbf{R}^{+}$, $\mu_{\max}(0)=\max\{\|\mu_{1}(0)\|,\ldots,\|\mu_{n}(0)\|\}$.
@@
$\mathcal{L}(\bm{\mu}(t+1))\leq\mathcal{L}(\bm{\mu}(t))-\frac{\eta}{2}\|\nabla\mathcal{L}(\bm{\mu}(t))\|^{2}\leq\mathcal{L}(\bm{\mu}(t))-\Omega\left(\frac{\eta\exp\left(-16n\mu_{\max}^{2}(0)\right)\pi_{\min}^{4}}{n^{2}d^{2}(1+\mu_{\max}(0){\sqrt{dn}})^{4}}\right)\mathcal{L}^{3}(\bm{\mu}(t)).$ (35)
@@
The first term $I_{1}$ can be bounded by Lemma 12 as
$I_{1}\geq\eta\Omega\left(\frac{\exp\left(-8\mu_{\max}^{2}(t)\right)\pi_{\min}^{2}}{d(1+\mu_{\max}(t){\sqrt{d}})^{2}}\right)\mu_{\max}^{4}(t)\geq\eta\Omega\left(\frac{\exp\left(-8n\mu_{\max}^{2}(0)\right)\pi_{\min}^{2}}{n^{2}d(1+\mu_{\max}(0){\sqrt{nd}})^{2}}\right)U^{2}(\bm{\mu}(t)).$ (37)
@@
$U(\bm{\mu}(t+1))\leq U(\bm{\mu}(t))-\eta\Omega\left(\frac{\exp\left(-8n\mu_{\max}^{2}(0)\right)\pi_{\min}^{2}}{n^{2}d(1+\mu_{\max}(0){\sqrt{nd}})^{2}}\right)U^{2}(\bm{\mu}(t))+\eta^{2}U(\bm{\mu}(t)).$
@@
a). If $\frac{n}{2}\mu_{\max}^{2}(0)\leq U(\bm{\mu}(t))\leq n\mu_{\max}^{2}(0)$, then
$\begin{split}&\quad U(\bm{\mu}(t+1))\leq U(\bm{\mu}(t))-\eta U(\bm{\mu}(t))\left(\Omega\left(\frac{\exp\left(-8n\mu_{\max}^{2}(0)\right)\pi_{\min}^{2}}{n^{2}d(1+\mu_{\max}(0){\sqrt{nd}})^{2}}\right)U(\bm{\mu}(t))-\eta\right)\\ &\leq U(\bm{\mu}(t))-\eta U(\bm{\mu}(t))\left(\Omega\left(\frac{\exp\left(-8n\mu_{\max}^{2}(0)\right)\pi_{\min}^{2}}{n^{2}d(1+\mu_{\max}(0){\sqrt{nd}})^{2}}\right)\frac{n}{2}\mu_{\max}^{2}(0)-\eta\right)\leq U(\bm{\mu}(t))\leq n\mu_{\max}^{2}(0),\end{split}$
note that we used $\eta\leq O\left(\frac{\exp\left(-8n\mu_{\max}^{2}(0)\right)\pi_{\min}^{2}}{n^{2}d(1+\mu_{\max}(0){\sqrt{nd}})^{2}}\right)\frac{n}{2}\mu_{\max}^{2}(0)$.
b). If $\frac{n}{2}\mu_{\max}^{2}(0)>U(\bm{\mu}(t))$, then $U(\bm{\mu}(t+1))\leq(1+\eta^{2})U(\bm{\mu}(t))\leq 2U(\bm{\mu}(t))\leq n\mu_{\max}^{2}(0)$.
@@
For any $\bm{\mu}$ satisfying $\|\mu_{1}\|\leq\sqrt{d},\|\mu_{2}\|,\|\mu_{3}\|,\ldots,\|\mu_{n}\|\geq 10\sqrt{d}$, the gradient of $\mathcal{L}$ at $\bm{\mu}$ can be upper bounded as
$\|\nabla_{\mu_{i}}\mathcal{L}(\bm{\mu})\|\leq 2\|\mu_{1}\|+2\exp(-d)\sum_{i\neq 1}\|\mu_{i}\|,\forall i\in[n].$

### v2

@@ 4.3 Proof ideas for Theorem 7
For any $\bm{\mu}$ satisfying $\|\mu_{1}\|,\|\mu_{2}\|\geq 10\sqrt{d},\|\mu_{3}\|,\ldots,\|\mu_{n}\|\leq\sqrt{d}$, the gradient of $\mathcal{L}$ at $\bm{\mu}$ can be upper bounded as
$\|\nabla_{\mu_{i}}\mathcal{L}(\bm{\mu})\|\leq 2(\|\mu_{3}\|+\cdots+\|\mu_{n}\|)+2\exp(-d)(\|\mu_{1}\|+\|\mu_{2}\|),\forall i\in[n].$
@@ 5 Experiments
Figure 1: Left: Sublinear convergence of the likelihood loss $\mathcal{L}$. Middle: Sublinear convergence of the parametric distance $\sum_{i\in[n]}\pi_{i}\|\mu_{i}-\mu^{*}\|^{2}$ between student GMM and the ground truth. Right: Impact of different mixing weights on the convergence speed.
Figure 2: Left: Gradient norm $\|\nabla\mathcal{L}(\bm{\mu}(0))\|$ in the counter-example in Theorem 7 decreases exponentially fast w.r.t. dimension $d$. Right: The statistical error (blue line) approximately scales as $\sim n^{-1/4}$ with sample size $n$.
In this section we experimentally explore the behavior of gradient EM on GMMs.
Convergence rates. We choose the experimental setting of $d=5,\eta=0.7$. We use $n=2,5,10$ Gaussian mixtures to learn data generated from one single ground truth Gaussian distribution $\mathcal{N}(\mu^{*},I_{d})$, respectively. Since a closed form expression of the population gradient is intractable, we approximate the gradient step via Monte Carlo method, with sample size $3.5\times 10^{5}$. The mixing weights of student GMM are randomly sampled from a standard Dirichlet distribution and set as fixed during gradient EM update. The covariances of all component Gaussians are set as the identity matrix. We recorded the convergence of likelihood function $\mathcal{L}$ (estimated also by Monte Carlo method on fresh samples each iteration) and parametric distance $\sum_{i\in[n]}\pi_{i}\|\mu_{i}-\mu^{*}\|^{2}$ along gradient EM trajectory. The results are reported in Figure 1 (left and middle panel). Both the likelihood $\mathcal{L}$ and the parametric distance converges sub-linearly.
Weight configurations. We train $3$-component GMM with $3$-different weight configurations and report $4$ runs each configuration in Figure 1 (right). Blue: $(\frac{1}{3},\frac{1}{3},\frac{1}{3})$. Orange: $(\frac{1}{6},\frac{1}{3},\frac{1}{2})$, Green: $(\frac{1}{20},\frac{1}{5},\frac{3}{4})$. More evenly distributed weights result in faster convergence.
Initialization geometry. We empirically study the bad initialization point $\bm{\mu}(0)$ described in Theorem 7  [^2] To prevent numerical underflow issues, we change the constant $12$ in $\bm{\mu}(0)$ to $2$. by plotting the gradient norm at $\bm{\mu}(0)$ w.r.t. different dimension $d$ in Figure 2 (left). As theoretically analyzed, the gradient norm $\|\nabla\mathcal{L}(\bm{\mu}(0))\|$ at $\bm{\mu}(0)$ decreases exponentially in dimension $d$.
Statistical rates. The statistical rate for EM/gradient EM is another interesting research problem, which we observe empirically in Figure 2 (right). We run gradient EM on $5$-component GMM with equal weights. x-axis: number of training samples, y-axis: parametric error after convergence. For each sample size, we run $50$ times and report the average. The statistical errors are reported in the blue line. The red line (function $\Theta(n^{-1/4})$) and green line (linear regression output fitting blue points) are references. The trajectory approximately follows the law of $accuracy\propto n^{-1/4}$. While [Wu and Zhou, 2019] rigorously proves the asymptotic statistical rate of $\tilde{O}(n^{-1/4})$ for the special case of 2-GMMs, our experiments imply that the same rate might also apply to the general case of multi-component GMMs.
@@ Appendix A Missing Proofs and Auxiliary lemmas
Recall that $U=\sum_{i\in[n]}\|\mu_{i}\|^{2}$. For any fixed $x\in\mathbf{R}^{d},x\neq 0$ and any $\bm{\mu}$ we have
$\int_{t=-1}^{1}\psi_{i}(tx|\bm{\mu})\psi_{j}(tx|\bm{\mu})\mathrm{d}t\geq\frac{1}{2\mu_{\max}\|x\|}\pi_{i}\pi_{j}\exp\left(-4U\right)\left(1-\exp\left(-4\mu_{\max}\|x\|\right)\right).$
@@
$\begin{split}\int_{t=-1}^{1}\psi_{i}(tx)\psi_{j}(tx)\mathrm{d}t&\geq\int_{t=-1}^{1}\pi_{i}\pi_{j}\exp\left(-4\mu_{\max}(\|tx\|+\mu_{\max})\right)\mathrm{d}t\\ &=\pi_{i}\pi_{j}\exp\left(-4\mu_{\max}^{2}\right)\cdot 2\int_{t=0}^{1}\exp\left(-4\mu_{\max}\|x\|t\right)\mathrm{d}t\\ &\geq\frac{1}{2\mu_{\max}\|x\|}\pi_{i}\pi_{j}\exp\left(-4U\right)\left(1-\exp\left(-4\mu_{\max}\|x\|\right)\right).\end{split}$ (13)
@@ B.1 Proofs for global convergence analysis
$\left\langle\nabla_{\bm{\mu}}Q(\bm{\mu}|\bm{\mu}),\bm{\mu}\right\rangle=\mathbf{E}_{x}[\|\tilde{\bm{\psi}}_{\bm{\mu}}(x)\|^{2}]\geq\Omega\left(\frac{\exp\left(-8U\right)\pi_{\min}^{2}}{d(1+\mu_{\max}{\sqrt{d}})^{2}}\mu_{\max}^{4}\right).$
@@
$\begin{split}\mathbf{E}_{x}\left[\|\tilde{\bm{\psi}}_{\bm{\mu}}(x)\|^{2}\right]&\geq\frac{\exp\left(-8U\right)}{40000d(1+2\mu_{\max}{\sqrt{d}})^{2}}\left(\sum_{i,j\in[n]}\pi_{i}\pi_{j}\|\mu_{i}-\mu_{j}\|^{2}\right)^{2}\\ &\geq\frac{\exp\left(-8U\right)}{40000d(1+2\mu_{\max}{\sqrt{d}})^{2}}\left(\frac{\pi_{\min}}{8}\mu_{\max}^{2}\right)^{2}\\ &=\frac{\exp\left(-8U\right)\pi_{\min}^{2}}{2560000d(1+2\mu_{\max}{\sqrt{d}})^{2}}\mu_{\max}^{4}.\end{split}$
Case2. For $\forall k\in[n]$, $\|\mu_{{i_{\max}}}-\mu_{k}\|<\frac{\mu_{\max}}{2}$. Then by Lemma 20 we have $\mathbf{E}_{x}\left[\|\tilde{\bm{\psi}}_{\bm{\mu}}(x)\|^{2}\right]\geq\frac{1}{4}\mu_{\max}^{2}\geq\Omega(\exp(-8\mu_{\max}^{2})\mu_{\max}^{4})\geq\Omega(\exp(-8U)\mu_{\max}^{4})\geq\Omega\left(\frac{\exp\left(-8U\right)\pi_{\min}^{2}}{d(1+\mu_{\max}{\sqrt{d}})^{2}}\mu_{\max}^{4}\right),$ (since $e^{-x}x\leq 1,\forall x$). ∎
@@
$\mathbf{E}_{x}\left[\|\tilde{\bm{\psi}}_{\bm{\mu}}(x)\|^{2}\right]\geq\frac{\exp\left(-8U\right)}{40000d(1+2\mu_{\max}{\sqrt{d}})^{2}}\left(\sum_{i,j\in[n]}\pi_{i}\pi_{j}\|\mu_{i}-\mu_{j}\|^{2}\right)^{2}.$
@@
$\begin{split}&\quad\mathbf{E}_{x}\left[\|\tilde{\bm{\psi}}_{\bm{\mu}}(x)\|^{2}\right]\\ &\geq\frac{1}{16}\mathbf{E}_{x}\left[\left(\int_{t=-1}^{1}\|x\|\sum_{i,j\in[n]}\psi_{i}(tx)\psi_{j}(tx)\langle\mu_{i}-\mu_{j},\overline{x}\rangle^{2}\mathrm{d}t\right)^{2}\right]\\ &=\frac{1}{16}\mathbf{E}_{x}\left[\left(\|x\|\sum_{i,j\in[n]}\langle\mu_{i}-\mu_{j},\overline{x}\rangle^{2}\int_{t=-1}^{1}\psi_{i}(tx)\psi_{j}(tx)\mathrm{d}t\right)^{2}\right]\\ &\geq\frac{1}{16}\mathbf{E}_{x}\left[\left(\|x\|\sum_{i,j\in[n]}\langle\mu_{i}-\mu_{j},\overline{x}\rangle^{2}\frac{1}{2\mu_{\max}\|x\|}\pi_{i}\pi_{j}\exp\left(-4U\right)\left(1-\exp\left(-4\mu_{\max}\|x\|\right)\right)\right)^{2}\right]\\ &=\frac{\exp\left(-8U\right)}{64}\mathbf{E}_{x}\left[\left(\sum_{i,j\in[n]}\pi_{i}\pi_{j}\langle\mu_{i}-\mu_{j},\overline{x}\rangle^{2}\frac{1-\exp\left(-4\mu_{\max}\|x\|\right)}{\mu_{\max}}\right)^{2}\right]\\ &\geq\frac{\exp\left(-8U\right)}{64}\left(\sum_{i,j\in[n]}\pi_{i}\pi_{j}\mathbf{E}_{x}\left[\langle\mu_{i}-\mu_{j},\overline{x}\rangle^{2}\frac{1-\exp\left(-4\mu_{\max}\|x\|\right)}{\mu_{\max}}\right]\right)^{2}\end{split}$ (24)
@@
$\begin{split}\mathbf{E}_{x}\left[\|\tilde{\bm{\psi}}_{\bm{\mu}}(x)\|^{2}\right]&\geq\frac{\exp\left(-8U\right)}{64}\left(\sum_{i,j\in[n]}\pi_{i}\pi_{j}\mathbf{E}_{x}\left[\langle\mu_{i}-\mu_{j},\overline{x}\rangle^{2}\frac{1-\exp\left(-4\mu_{\max}\|x\|\right)}{\mu_{\max}}\right]\right)^{2}\\ &\geq\frac{\exp\left(-8U\right)}{64}\left(\sum_{i,j\in[n]}\pi_{i}\pi_{j}\frac{1-\exp\left(-2\mu_{\max}{\sqrt{d}}\right)}{50d\mu_{\max}}\|\mu_{i}-\mu_{j}\|^{2}\right)^{2}\\ &\geq\frac{\exp\left(-8U\right)}{64}\left(\sum_{i,j\in[n]}\pi_{i}\pi_{j}\frac{1-\frac{1}{1+2\mu_{\max}{\sqrt{d}}}}{50d\mu_{\max}}\|\mu_{i}-\mu_{j}\|^{2}\right)^{2}\\ &=\frac{\exp\left(-8U\right)}{40000d(1+2\mu_{\max}{\sqrt{d}})^{2}}\left(\sum_{i,j\in[n]}\pi_{i}\pi_{j}\|\mu_{i}-\mu_{j}\|^{2}\right)^{2}\end{split}$ (28)
@@
Consider training a student $n$-component GMM initialized from $\bm{\mu}(0)=(\mu_{1}(0)^{\top},\ldots,\mu_{n}(0)^{\top})^{\top}$ to learn a single-component ground truth GMM $\mathcal{N}(0,I_{d})$ with population gradient EM algorithm. If the step size satisfies $\eta\leq O\left(\frac{\exp\left(-8U(0)\right)\pi_{\min}^{2}}{n^{2}d^{2}(\frac{1}{\mu_{\max}(0)}+\mu_{\max}(0))^{2}}\right)$, then gradient EM converges globally with rate
@@
where $\gamma=\Omega\left(\frac{\eta\exp\left(-16U(0)\right)\pi_{\min}^{4}}{n^{2}d^{2}(1+\mu_{\max}(0){\sqrt{dn}})^{4}}\right)\in\mathbf{R}^{+}$. Recall that $\mu_{\max}(0)=\max\{\|\mu_{1}(0)\|,\ldots,\|\mu_{n}(0)\|\}$ and $U(0)=\sum_{i\in[n]}\|\mu_{i}(0)\|^{2}$ are two initialization constants.
@@
$\mathcal{L}(\bm{\mu}(t+1))\leq\mathcal{L}(\bm{\mu}(t))-\frac{\eta}{2}\|\nabla\mathcal{L}(\bm{\mu}(t))\|^{2}\leq\mathcal{L}(\bm{\mu}(t))-\Omega\left(\frac{\eta\exp\left(-16U(0)\right)\pi_{\min}^{4}}{n^{2}d^{2}(1+\mu_{\max}(0){\sqrt{dn}})^{4}}\right)\mathcal{L}^{3}(\bm{\mu}(t)).$ (35)
@@
By induction hypothesis, the first term $I_{1}$ can be bounded by Lemma 12 as
$I_{1}\geq\eta\Omega\left(\frac{\exp\left(-8U(t)\right)\pi_{\min}^{2}}{d(1+\mu_{\max}(t){\sqrt{d}})^{2}}\right)\mu_{\max}^{4}(t)\geq\eta\Omega\left(\frac{\exp\left(-8U(0)\right)\pi_{\min}^{2}}{n^{2}d(1+\mu_{\max}(0){\sqrt{nd}})^{2}}\right)U^{2}(\bm{\mu}(t)).$ (37)
@@
$U(\bm{\mu}(t+1))\leq U(\bm{\mu}(t))-\eta\Omega\left(\frac{\exp\left(-8U(0)\right)\pi_{\min}^{2}}{n^{2}d(1+\mu_{\max}(0){\sqrt{nd}})^{2}}\right)U^{2}(\bm{\mu}(t))+\eta^{2}U(\bm{\mu}(t)).$
@@
a). If $\frac{U(0)}{2}\leq U(\bm{\mu}(t))\leq U(0)$, then
$\begin{split}&\quad U(\bm{\mu}(t+1))\leq U(\bm{\mu}(t))-\eta U(\bm{\mu}(t))\left(\Omega\left(\frac{\exp\left(-8U(0)\right)\pi_{\min}^{2}}{n^{2}d(1+\mu_{\max}(0){\sqrt{nd}})^{2}}\right)U(\bm{\mu}(t))-\eta\right)\\ &\leq U(\bm{\mu}(t))-\eta U(\bm{\mu}(t))\left(\Omega\left(\frac{\exp\left(-8U(0)\right)\pi_{\min}^{2}}{n^{2}d(1+\mu_{\max}(0){\sqrt{nd}})^{2}}\right)\frac{n}{2}\mu_{\max}^{2}(0)-\eta\right)\leq U(\bm{\mu}(t))\leq n\mu_{\max}^{2}(0),\end{split}$
note that we used $\eta\leq O\left(\frac{\exp\left(-8U(0)\right)\pi_{\min}^{2}}{n^{2}d(1+\mu_{\max}(0){\sqrt{nd}})^{2}}\right)\frac{n}{2}\mu_{\max}^{2}(0)$.
b). If $U(\bm{\mu}(t))<\frac{1}{2}U(0)$, then $U(\bm{\mu}(t+1))\leq(1+\eta^{2})U(\bm{\mu}(t))\leq 2U(\bm{\mu}(t))\leq U(0)$.
@@
For any $\bm{\mu}$ satisfying $\|\mu_{1}\|,\|\mu_{2}\|\geq 10\sqrt{d},\|\mu_{3}\|,\ldots,\|\mu_{n}\|\leq\sqrt{d}$, the gradient of $\mathcal{L}$ at $\bm{\mu}$ can be upper bounded as
$\|\nabla_{\mu_{i}}\mathcal{L}(\bm{\mu})\|\leq 2(\|\mu_{3}\|+\cdots+\|\mu_{n}\|)+2\exp(-d)(\|\mu_{1}\|+\|\mu_{2}\|),\forall i\in[n].$